# Chronos: Learning the Language of Time Series

### v2

@@ 3.1 Time Series Tokenization
The scale of time series can differ significantly even within a single dataset. This poses optimization challenges for deep learning models. Therefore, individual time series are normalized to facilitate better optimization. In the case of Chronos, the goal of normalization is to map the time series values into a suitable range for quantization. A common normalization technique involves applying an affine transformation to the time series, i.e., $\tilde{x}_{i}={(x_{i}-m)}/{s}$. Several popular normalization schemes, such as mean scaling, standard scaling and min-max scaling, can be obtained by appropriately choosing $m$ and $s$. We opt for mean scaling, a method that has proven effective in deep learning models commonly used for practical time series applications (Salinas et al., 2020), but other approaches are viable and only require minimal changes. Mean scaling normalizes individual entries of the time series by the mean of the absolute values in the historical context. Specifically, this involves setting $m=0$ and $s=\frac{1}{C}\sum_{i=1}^{C}|x_{i}|$.
@@
respectively. The positioning of bin centers and edges can either be data-dependent or uniform (Rabanser et al., 2020). Quantile binning, a type of data-dependent binning, exploits the cumulative distribution function (CDF) of the training datapoints to construct bins such that approximately equal number of datapoints are assigned to each bin. In contrast, uniform binning selects bin centers uniformly within some interval $[l,r]$. Since the distribution of values for unseen downstream datasets can differ significantly from the training distribution, we opt for uniform binning in our experiments, but other quantization techniques can be used. We refer the reader to Rabanser et al. (2020) for a detailed discussion on quantization schemes for time series. A potential limitation of this approach is that the prediction range is restricted between $[c_{1},c_{B}],$ making it theoretically infeasible to model time series with a strong trend. We explore this further in a practical setting in Section 5.7.
@@ 3.2 Objective Function
Note that the categorical cross entropy loss (Eq. 2) is not a distance-aware objective function, i.e., it does not explicitly recognize that bin $i$ is closer to bin $i+1$ than to $i+2$. Instead, the model is expected to associate nearby bins together, based on the distribution of bin indices in the training dataset. In other words, Chronos performs regression via classification (Torgo & Gama, 1997). This is unlike typical probabilistic time series forecasting models, which either use parametric continuous distributions such as Gaussian and Student’s-t (Salinas et al., 2020) or perform quantile regression (Wen et al., 2017; Lim et al., 2021).
@@ 4.1 TSMixup: Time Series Mixup
Mixup (Zhang et al., 2017) is a data augmentation scheme proposed in the context of image classification. It generates convex combinations of random image pairs and their labels from the training dataset, which alleviates issues such as memorization and overfitting in deep learning models. Existing works (Carmona et al., 2021; Zhou et al., 2023b) have extended Mixup to the time series domain. Building upon these works, we propose TSMixup, which generalizes the idea of Mixup to more than two datapoints. Concretely, TSMixup randomly samples $k\sim\mathcal{U}\{1,K\}$ time series of a specific length, $l\sim\mathcal{U}\{l_{\min},l_{\max}\}$, from the training datasets, scales them, and takes their convex combination,
@@
where $\tilde{{\bm{x}}}^{(i)}_{1:l}$ denotes the $i$-th scaled time series. The combination weights, $[\lambda_{1},\dots,\lambda_{k}]$, are sampled from a symmetric Dirichlet distribution, $\mathrm{Dir}(\alpha)$. The complete pseudocode of TSMixup can be found in Algorithm 1 in Appendix A. Intuitively, TSMixup enhances the diversity of data by combining patterns from different time series. Figure 2 shows example augmentations generated by TSMixup and illustrates how different patterns are mixed.
@@ 5 Experiments
Benchmark I 15 97,272 pretraining and in-domain evaluation Naive, SeasonalNaive, AutoETS, AutoTheta, AutoARIMA, DeepAR, TFT, PatchTST, DLinear, WaveNet, N-BEATS, N-HiTS, GPT4TS, Lag-Llama, Moirai-1.0-R
@@ 5.1 Datasets
To train and evaluate Chronos models, we collected a wide variety of publicly available datasets spanning various application domains including energy, transport, healthcare, retail, web, weather, finance, and with sampling frequencies ranging from 5 minutes up to yearly. The complete list of datasets, together with their respective sources and additional details, is given in Appendix B. In total, our dataset collection comprises 55 datasets from multiple sources, including the Monash Time Series Forecasting Repository (Godahewa et al., 2021), the M-competitions (Makridakis et al., 1979; Makridakis & Hibon, 2000; Makridakis et al., 2020; 2022), and public domain datasets from Kaggle.
@@ 5.2 Training Corpus and Protocols
We trained T5 models of 4 sizes, [^1] Our code and model checkpoints are available at https://github.com/amazon-science/chronos-forecasting. namely, Mini (20M), Small (46M), Base (200M) and Large (710M), and the GPT-2 base model (90M), on 10M TSMixup augmentations (see Section 4.1) generated from the 28 training datasets, with $K=3$ in Algorithm 1, and 1M synthetic time series generated using Gaussian processes (see Section 4.2). Note that with this setup, original time series are adequately represented since they are included in the TSMixup augmentations with probability $1/3$. We sampled time series from the augmentations and synthetic data in the ratio 9:1 during training. Each model is trained with an effective batch size of 256 sequences, using distributed data parallelism and gradient accumulation, whenever necessary. These sequences are constructed by slicing random windows from the time series, and then scaling and quantizing them into equal-sized bins within the interval $[l{=}-15$, $r{=}15]$, as described in Section 3.1. The context length of the sequences was set to 512, the default for T5 models, and the prediction length is set to 64, a value greater than the prediction lengths of all tasks we consider in our evaluation.
@@ 5.3 Baselines
We assessed the performance of Chronos models against a variety of time series forecasting baselines. From statistical forecasting literature (Hyndman & Athanasopoulos, 2018), we included Naive, Seasonal Naive, AutoETS, AutoARIMA (Hyndman et al., 2008) and AutoTheta (Assimakopoulos & Nikolopoulos, 2000). Additionally, we compared against several neural forecasting baselines, including WaveNet (Oord et al., 2016), DeepAR (Salinas et al., 2020), N-BEATS (Oreshkin et al., 2020), TFT (Lim et al., 2021), DLinear (Zeng et al., 2023), PatchTST (Nie et al., 2023), N-HiTS (Challu et al., 2023), and GPT4TS (Zhou et al., 2023a). Furthermore, from the recently proposed pretrained time series models, we included the ones with publicly available weights: Lag-Llama (Rasul et al., 2023) and Moirai-1.0-R (Woo et al., 2024). On Benchmark II (i.e., zero-shot datasets for Chronos models), we also evaluated against two zero-shot methods: ForecastPFN (Dooley et al., 2023) which is a transformer model pretrained only on synthetic time series data and LLMTime (Gruver et al., 2023) which uses LLMs for zero-shot forecasting.
@@ 5.4 Evaluation Metrics
Since the magnitude of the evaluation metrics can vary across datasets, we adopt a different approach to aggregate scores than naive averaging. For each dataset, we compute the relative score of each model as the model’s score divided by the score of a baseline model (here, Seasonal Naive). The relative scores are aggregated across all datasets using the geometric mean. The choice of the geometric mean is deliberate —  Fleming & Wallace (1986) show that the arithmetic mean can yield misleading conclusions in this context, and the geometric mean is provably the only meaningful way to aggregate such relative scores. Furthermore, the geometric mean is also not sensitive to the choice of the baseline, and the model ordering stays intact if another baseline is selected instead. We used Seasonal Naive due to its simplicity and popularity as a forecasting baseline. For models that failed or could not finish evaluation within the allotted time on certain datasets, we use a relative score of 1, i.e., the baseline relative score, when aggregating the results. We assign equal weights to all tasks during aggregation, reflecting real-world scenarios where datasets may have different numbers of time series, frequencies, history and prediction lengths.
@@ 5.5.1 Benchmark I: In-domain Results
Figure 4: Performance of different models on Benchmark I, comprising 15 datasets also included in the training data of Chronos models. This benchmark showcases the in-domain performance of Chronos models against local statistical models, which fit parameters individually for each time series, task-specific models that train a separate model for each task, and pretrained models trained on a large corpus of time series data. Pretrained Models (Other) indicates that the in-domain setting does not apply to these models as some datasets in Benchmark I were not part of their training corpus and (or) they were trained on the test sets of some datasets in Benchmark I. The probabilistic (WQL) and point (MASE) forecasting metrics are normalized using the scores of the Seasonal Naive baseline and aggregated through a geometric mean to obtain the aggregated relative WQL and MASE, respectively. Results for Chronos and task-specific models (except GPT4TS) have been averaged over 3 random seeds. Models producing point-forecasts (GPT4TS) are only compared based on MASE.
@@
The smaller Chronos-T5 models (Mini and Small) and Chronos-GPT2 also perform better than the majority of baselines, with the exception of PatchTST. Between the two baseline pretrained models studied in this experiment, Moirai-1.0-R clearly outperforms Lag-Llama. Notably, the best Moirai-1.0-R model (Large, 311M) is still outperformed by the smallest Chronos-T5 model (Mini, 20M) even though Moirai-1.0-R models were trained on a significantly larger corpus of time series data. Task-specific deep learning models, trained across multiple time series for a specific task, perform better than local statistical models that fit parameters for each time series. Interestingly, the Seasonal Naive baseline performs competitively against other local models on this benchmark, suggesting that the datasets in this benchmark exhibit strong seasonal patterns. This is unsurprising since a majority of these datasets belong to domains such as energy and transport that tend to be highly seasonal in nature. The raw WQL and MASE values for individual datasets summarized in Figure 4 can be found in Tables 6 and 7 in Appendix E.
@@ 5.5.2 Benchmark II: Zero-shot Results
Figure 5: Performance of different models on Benchmark II, comprising 27 datasets not seen by Chronos models during training. This benchmark provides insights into the zero-shot performance of Chronos models against local statistical models, which fit parameters individually for each time series, task-specific models trained on each task, and pretrained models trained on a large corpus of time series data. Pretrained Models (Other) indicates that the zero-shot setting does not apply to these models as they were trained on some datasets in Benchmark II. The probabilistic (WQL) and point (MASE) forecasting metrics were normalized using the scores of the Seasonal Naive baseline and aggregated through a geometric mean to obtain the aggregated relative WQL and MASE, respectively. Results for Chronos and task-specific models (except GPT4TS) have been averaged over 3 random seeds. Models producing point-forecasts (GPT4TS and ForecastPFN) are only compared based on MASE.
@@
Nevertheless, despite never having seen these datasets during training, Chronos models significantly outperform local statistical models. On probabilistic forecasting (aggregate relative WQL), Chronos models achieve the 2nd and 3rd spots, performing better than most task-specific models that have been trained on these tasks. Chronos-T5 (Large) places 3rd in terms of the point forecasting performance, narrowly losing the 2nd spot to N-HiTS. Chronos models also significantly outperform other pretrained models such as Moirai-1.0-R, Lag-Llama, LLMTime, and ForecastPFN, and even GPT4TS, which fine-tunes a pretrained GPT-2 model on each dataset. Moirai-1.0-R obtains the best performance after Chronos although the evaluation setup may have been advantageous for Moirai-1.0-R as many datasets in Benchmark II were part of its pretraining corpus. The raw WQL and MASE values for individual datasets summarized in Figure 5 can be found in Tables 8 and 9 in Appendix E.
@@ 5.6 Analysis of Hyperparameters
Figure 7: Model size. (a) Training loss curves of Chronos models of different sizes. (b) In-domain and zero-shot performance of Chronos models varying over model size.
@@
Figure 9: Comparison of the in-domain and zero-shot performance of models initialized with language model weights (marked as star) and three randomly initialized models (marked as circles) across different model sizes.
@@
We studied the effect of the context length on downstream performance by training Chronos-T5 (Small, 46M) models with four distinct context lengths. Figure 11b shows how the performance varies with increasing context length. We observe improvements on both in-domain and zero-shot metrics as context length increases, showing that a longer context helps the models to forecast better. However, this analysis may be limited due to our zero-shot evaluation setup, wherein the majority of datasets in the benchmark have low frequencies and time series shorter than 1000 steps. Hence, further evaluation is required to conclusively study the impact of longer context lengths. We posit that high-frequency datasets may benefit from a longer context, which may be necessary to correctly capture the long-term seasonal patterns.
@@
The vocabulary size governs the precision with which the model can process the scaled time series. To explore its impact on performance, we trained Chronos-T5 (Small, 46M) models with varying vocabulary sizes. Figure 11c shows consistent improvements in the point forecasting metric (MASE) as the vocabulary size increases. In contrast, the WQL initially improves but deteriorates for larger vocabulary sizes. We hypothesize that this behavior is an artifact of the chosen metrics. The MASE, which is invariant to the scale of individual series, is closely aligned to our training loss, which is also invariant to scale. Hence, MASE exhibits an improvement with increased precision, just as one expects for the training loss. Conversely, WQL, a scale-dependent metric, does not correlate closely with the training loss and behaves less predictably as precision increases. See Appendix D for a discussion on the properties of these metrics.

### v3

@@ 3.1 Time Series Tokenization
The scale of time series can differ significantly even within a single dataset. This poses optimization challenges for deep learning models. Therefore, individual time series are normalized to facilitate better optimization. In the case of Chronos, the goal of normalization is to map the time series values into a suitable range for quantization. A common normalization technique involves applying an affine transformation to the time series, i.e., $\tilde{x}_{i}={(x_{i}-m)}/{s}$. Several popular normalization schemes, such as mean scaling, standard scaling and min-max scaling, can be obtained by appropriately choosing $m$ and $s$. We opt for mean scaling, a method that has proven effective in deep learning models commonly used for practical time series applications (Salinas et al., 2020; Rabanser et al., 2020), but other approaches are viable and only require minimal changes. An attractive feature of mean scaling is that it preserves zero values in the time series, which are often semantically meaningful, such as zero sales for a product or zero solar energy generation at night. Mean scaling normalizes individual entries of the time series by the mean of the absolute values in the historical context. Specifically, this involves setting $m=0$ and $s=\frac{1}{C}\sum_{i=1}^{C}|x_{i}|$.
@@
respectively. The positioning of bin centers and edges can either be data-dependent or uniform (Rabanser et al., 2020). Quantile binning, a type of data-dependent binning, exploits the cumulative distribution function (CDF) of the training datapoints to construct bins such that approximately equal number of datapoints are assigned to each bin. In contrast, uniform binning selects uniformly-spaced bin centers within the interval $[c_{1},c_{B}]$ and the bin edges fall mid-way between the successive bin centers, i.e., $b_{i}=\frac{c_{i}+c_{i+1}}{2}$ for $i\in\{1,\dots,B-1\}$. Since the distribution of values for unseen downstream datasets can differ significantly from the training distribution, we opt for uniform binning in our experiments, but other quantization techniques can be used. We refer the reader to Rabanser et al. (2020) for a detailed discussion on quantization schemes for time series. A potential limitation of this approach is that the prediction range is restricted between $[c_{1},c_{B}],$ making it theoretically infeasible to model time series with a strong trend. We explore this further in a practical setting in Section 5.7.
@@ 3.2 Objective Function
Note that the categorical cross entropy loss (Eq. 2) is not a distance-aware objective function, i.e., it does not explicitly recognize that bin $i$ is closer to bin $i+1$ than to $i+2$. Instead, the model is expected to associate nearby bins together, based on the distribution of bin indices in the training dataset. In other words, Chronos performs regression via classification (Torgo & Gama, 1997; Stewart et al., 2023). This is unlike typical probabilistic time series forecasting models, which either use parametric continuous distributions such as Gaussian and Student’s-t (Salinas et al., 2020) or perform quantile regression (Wen et al., 2017; Lim et al., 2021).
@@
Arguably, modeling the output as an ordinal variable would be more appropriate, since the output domain is obtained by discretizing the real line. In fact, regression models for ordinal variables have been extensively studied in the literature (McCullagh, 1980; Winship & Mare, 1984), including for neural networks and transformer models (Cheng et al., 2008; Hu et al., 2021). Imposing the ordinal nature of the classes on top of the models, in similar ways to the mentioned literature, could be an interesting extension of this work.
@@ 4.1 TSMixup: Time Series Mixup
Mixup (Zhang et al., 2017) is a data augmentation scheme proposed in the context of image classification. It generates convex combinations of random image pairs and their labels from the training dataset, which alleviates issues such as memorization and overfitting in deep learning models. Existing works (Carmona et al., 2021; Zhou et al., 2023b) have extended Mixup to the time series domain.
@@
Building upon these works, we propose TSMixup, which generalizes the idea of Mixup to more than two datapoints. Concretely, TSMixup randomly samples $k\sim\mathcal{U}\{1,K\}$ time series of a specific length, $l\sim\mathcal{U}\{l_{\min},l_{\max}\}$, from the training datasets, scales them, and takes their convex combination,
@@
where $\tilde{{\bm{x}}}^{(i)}_{1:l}$ denotes the $i$-th scaled time series. The time series are scaled before mixing to ensure that time series with small and large values are given equal importance in the mixing process. The combination weights, $[\lambda_{1},\dots,\lambda_{k}]$, are sampled from a symmetric Dirichlet distribution, $\mathrm{Dir}(\alpha)$, parameterized by the scalar concentration parameter $\alpha$. The complete pseudocode of TSMixup can be found in Algorithm 1 in Appendix A. Intuitively, TSMixup enhances the diversity of data by combining patterns from different time series. Figure 2 shows example augmentations generated by TSMixup and illustrates how different patterns are mixed.
@@ 5 Experiments
Benchmark I 15 97,272 pretraining and in-domain evaluation Naive, SeasonalNaive, AutoETS, AutoTheta, SCUM, AutoARIMA, DeepAR, TFT, PatchTST, DLinear, WaveNet, N-BEATS, N-HiTS, GPT4TS, Lag-Llama, Moirai-1.0-R
@@ 5.1 Datasets
To train and evaluate Chronos models, we collected a wide variety of publicly available datasets spanning various application domains including energy, transport, healthcare, retail, web, weather, finance, and with sampling frequencies ranging from 5 minutes up to yearly. The complete list of datasets, together with their respective sources and additional details, is given in Appendix B. In total, our dataset collection comprises 55 datasets from multiple sources, including the Monash Time Series Forecasting Repository (Godahewa et al., 2021), the M-competitions (Makridakis et al., 1979; Makridakis & Hibon, 2000; Makridakis et al., 2020; 2022), and public domain datasets from Kaggle. [^1] The datasets used in our experiments are available at https://huggingface.co/datasets/autogluon/chronos_datasets.
@@ 5.2 Training Corpus and Protocols
We trained T5 models of 4 sizes, [^2] Our code and model checkpoints are available at https://github.com/amazon-science/chronos-forecasting. namely, Mini (20M), Small (46M), Base (200M) and Large (710M), and the GPT-2 base model (90M), on 10M TSMixup augmentations (see Section 4.1) generated from the 28 training datasets, with $K=3$ in Algorithm 1, and 1M synthetic time series generated using Gaussian processes (see Section 4.2). Note that with this setup, original time series are adequately represented since they are included in the TSMixup augmentations with probability $1/3$. We sampled time series from the augmentations and synthetic data in the ratio 9:1 during training. Each model is trained with an effective batch size of 256 sequences, using distributed data parallelism and gradient accumulation, whenever necessary. These sequences were constructed by slicing random windows from the time series, and then scaling and quantizing them into equal-sized bins within the interval $[c_{1}{=}-15$, $c_{B}{=}+15]$, as described in Section 3.1. We set the vocabulary size, ${\mathcal{V}}_{\mathrm{ts}}$, to 4096, including the special tokens (PAD and EOS). The context length of the sequences was set to 512, the default for T5 models, and the prediction length was set to 64, a value greater than the prediction lengths of all tasks we consider in our evaluation.
@@ 5.3 Baselines
We assessed the performance of Chronos models against a variety of time series forecasting baselines. From statistical forecasting literature (Hyndman & Athanasopoulos, 2018), we included Naive, Seasonal Naive, AutoETS, AutoARIMA (Hyndman et al., 2008), AutoTheta (Assimakopoulos & Nikolopoulos, 2000) and a strong ensemble (SCUM) of statistical models (Petropoulos & Svetunkov, 2020). Additionally, we compared against several neural forecasting baselines, including WaveNet (Oord et al., 2016), DeepAR (Salinas et al., 2020), N-BEATS (Oreshkin et al., 2020), TFT (Lim et al., 2021), DLinear (Zeng et al., 2023), PatchTST (Nie et al., 2023), N-HiTS (Challu et al., 2023), and GPT4TS (Zhou et al., 2023a). Furthermore, from the recently proposed pretrained time series models, we included the ones with publicly available weights: Lag-Llama (Rasul et al., 2023) and Moirai-1.0-R (Woo et al., 2024). On Benchmark II (i.e., zero-shot datasets for Chronos models), we also evaluated against two zero-shot methods: ForecastPFN (Dooley et al., 2023) which is a transformer model pretrained only on synthetic time series data and LLMTime (Gruver et al., 2023) which uses LLMs for zero-shot forecasting.
@@ 5.4 Evaluation Metrics
Since the magnitude of the evaluation metrics can vary across datasets, we adopt a different approach to aggregate scores than naive averaging. For each dataset, we compute the relative score of each model as the model’s score divided by the score of a baseline model (here, Seasonal Naive). The relative scores are aggregated across all datasets using the geometric mean. The choice of the geometric mean is deliberate —  Fleming & Wallace (1986) show that the arithmetic mean can yield misleading conclusions in this context, and the geometric mean is provably the only meaningful way to aggregate such relative scores. Furthermore, the geometric mean is also not sensitive to the choice of the baseline, and the model ordering stays intact if another baseline is selected instead. We used Seasonal Naive due to its simplicity and popularity as a forecasting baseline. For models that failed or could not finish evaluation within the allotted time on certain datasets, we used a relative score of 1, i.e., the baseline relative score, when aggregating the results. We assign equal weights to all tasks during aggregation, reflecting real-world scenarios where datasets may have different numbers of time series, frequencies, history and prediction lengths.
@@ 5.5.1 Benchmark I: In-domain Results
Figure 4: Performance of different models on Benchmark I, comprising 15 datasets also included in the training data of Chronos models. This benchmark showcases the in-domain performance of Chronos models against local statistical models, which fit parameters individually for each time series, task-specific models that train a separate model for each task, and pretrained models trained on a large corpus of time series data. Pretrained Models (Other) indicates that the in-domain setting does not apply to these models as they were trained on different corpora than Chronos. Specifically, this means that some datasets in Benchmark I were not part of their training corpus and (or) they were trained on the test sets of some datasets in Benchmark I. The probabilistic (WQL) and point (MASE) forecasting metrics (lower is better) are normalized using the scores of the Seasonal Naive baseline and aggregated through a geometric mean to obtain the aggregated relative WQL and MASE, respectively. Results for Chronos and task-specific models (except GPT4TS) have been averaged over 3 random seeds. Models producing point-forecasts (GPT4TS) are only compared based on MASE.
@@
The smaller Chronos-T5 models (Mini and Small) and Chronos-GPT2 also perform better than the majority of baselines. Between the two baseline pretrained models studied in this experiment, Moirai-1.0-R clearly outperforms Lag-Llama. Notably, the best Moirai-1.0-R model (Large, 311M) is still outperformed by the smallest Chronos-T5 model (Mini, 20M) even though Moirai-1.0-R models were trained on a significantly larger corpus of time series data. Task-specific deep learning models, trained across multiple time series for a specific task, perform better than local statistical models that fit parameters for each time series. Interestingly, the Seasonal Naive baseline performs competitively against other local models on this benchmark, suggesting that the datasets in this benchmark exhibit strong seasonal patterns. This is unsurprising since a majority of these datasets belong to domains such as energy and transport that tend to be highly seasonal in nature. The raw WQL and MASE values for individual datasets summarized in Figure 4 can be found in Tables 7 and 8 in Appendix E.
@@ 5.5.2 Benchmark II: Zero-shot Results
Figure 5: Performance of different models on Benchmark II, comprising 27 datasets not seen by Chronos models during training. This benchmark provides insights into the zero-shot performance of Chronos models against local statistical models, which fit parameters individually for each time series, task-specific models trained on each task, and pretrained models trained on a large corpus of time series data. Pretrained Models (Other) indicates that the zero-shot setting does not apply to these models as they were pretrained on some datasets in Benchmark II. The probabilistic (WQL) and point (MASE) forecasting metrics (lower is better) were normalized using the scores of the Seasonal Naive baseline and aggregated through a geometric mean to obtain the aggregated relative WQL and MASE, respectively. Results for Chronos and task-specific models (except GPT4TS) have been averaged over 3 random seeds. Models producing point-forecasts (GPT4TS and ForecastPFN) are only compared based on MASE.
@@
Nevertheless, despite never having seen these datasets during training, Chronos models significantly outperform standalone local statistical models. On probabilistic forecasting (aggregate relative WQL), Chronos models achieve the 2nd to 4th spots, performing better than most task-specific models that have been trained on these tasks. In terms of the point forecasting performance, Chronos-T5 (Large) places 2nd, surpassing most baselines, including the strong SCUM ensemble. Chronos models also significantly outperform other pretrained models such as Moirai-1.0-R, Lag-Llama, LLMTime, and ForecastPFN, and even GPT4TS, which fine-tunes a pretrained GPT-2 model on each dataset. Moirai-1.0-R obtains the best performance after Chronos, although the evaluation setup may have been advantageous for Moirai-1.0-R as many datasets in Benchmark II were part of its pretraining corpus. The raw WQL and MASE values for individual datasets summarized in Figure 5 can be found in Tables 9 and 10 in Appendix E.
@@ 5.6 Analysis of Hyperparameters
Figure 7: Model size. (a) Training loss curves of Chronos models of different sizes. (b) In-domain and zero-shot performance of Chronos models varying over model size (lower is better).
@@
Figure 9: Comparison of the in-domain and zero-shot performance (lower is better) of models initialized with language model weights (marked as star) and three randomly initialized models (marked as circles) across different model sizes.
@@
We studied the effect of the context length on downstream performance by training Chronos-T5 (Small, 46M) models with four distinct context lengths. Figure 11b shows how the performance varies with increasing context length. We observe improvements on both in-domain and zero-shot metrics as context length increases up to 1024, showing that a longer context helps the models to forecast better to a certain degree. However, increasing the context length further tends to saturate or worsen the performance, which may partly be due to a limitation of our evaluation setup: it does not include enough high-frequency datasets (>=>=> = 15 min). Hence, further evaluation is required to conclusively study the impact of longer context lengths. We posit that high-frequency datasets may benefit from a longer context, which may be necessary to correctly capture the long-term seasonal patterns.
@@
The vocabulary size governs the precision with which the model can process the scaled time series. To explore its impact on performance, we trained Chronos-T5 (Small, 46M) models with varying vocabulary sizes. Figure 11c shows modest improvements in the point forecasting metric (MASE) as the vocabulary size increases. In contrast, the WQL initially improves but deteriorates for larger vocabulary sizes. We hypothesize that this behavior is an artifact of the chosen metrics. The MASE, which is invariant to the scale of individual series, is closely aligned to our training loss, which is also invariant to scale. Hence, MASE exhibits an improvement with increased precision, just as one expects for the training loss. Conversely, WQL, a scale-dependent metric, does not correlate closely with the training loss and behaves less predictably as precision increases. See Appendix D for a discussion on the properties of these metrics. Beyond this experiment, we posit that selecting the vocabulary size in the context of a model like Chronos would pose a trade-off. A vocabulary that is too small would lead to poor forecasting accuracy due to large discretization errors; however, a large vocabulary would lead to the bins being too fine, potentially leading to generalization errors due to fewer datapoints falling into each bin.